OpenStudy (anonymous):

diethyl ether is produced from ethanol according to the following equation 2CH3CH2OH--->CH3CH2OCH2CH3+H2O What mass of ether will be produced if 221 g of ethanol reacts with a 66.5 percent yield??

OpenStudy (aaronq):

find the theoretical yield and multiply it by the percent yield (0.665).

OpenStudy (aaronq):

if you need finding the theoretical yield, convert the mass given of reactant to moles and follow this: Use the stoichiometric coefficients to find moles produced. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) -------------------------------------------------------- To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles.